OpenStudy (anonymous):

Find f(g(x)). What does your answer mean? g(x)=1/2(x)- (1/3) f(x)= (6x+2)/3

OpenStudy (unklerhaukus):

\[f(x)= \frac{(6x+2)}{3}\] \[g(x)=\frac12x-\frac13\] \[f(g(x))=f\circ g(x)=\frac{\left(6\left(\frac12x-\frac13\right)+2\right)}{3}\] \[=\frac{3x-2+2}{3}\]\[=\cdots\]